# Principal stratification with $U$-statistics under principal ignorability

### v4

@@ 4.4 Sensitivity analysis for monotonicity
To point identify PGCEs, a central assumption is monotonicity (Assumption 2) which rules out the $01$ stratum. This stratum may be relevant, for example, when comparing two active interventions with unknown survival benefits against non-mortality outcomes. We thus describe a sensitivity analysis strategy to evaluate the robustness of the proposed estimators in the event of departure from monotonicity. Following the approach in Ding and Lu (2016), we define the sensitivity function $\eta(X)$ as the ratio between the proportion of strata $01$ and $10$ conditional on the covariates, i.e., $\eta(X)=\mathbb{P}(S=01|X)/\mathbb{P}(S=10|X)$, where $\eta(X)\in[0,\infty)$; the monotonicity holds if $\eta(X)=0$, but $\eta(X)>0$ implies the existence of the $01$ stratum. The PCGEs in this scenario are thus $\tau_{h,\eta}^{s}$ for $s\in\{10,01,00,11\}$. To proceed, we first notice that, for $\eta(X)\neq 1$, the principal scores can be obtained as:
$e_{10,\eta}(X)=\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},~~e_{01,\eta}(X)=\frac{\eta(X)\{p_{1}(X)-p_{0}(X)\}}{1-\eta(X)},$ (15)
$e_{00,\eta}(X)=1-p_{0}(X)-\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},~~e_{11,\eta}(X)=p_{1}(X)-\frac{p_{1}(X)-p_{0}(X)}{1-\eta(X)},$
with $0\leq\eta(X)\leq 1-\{p_{1}(X)-p_{0}(X)\}/\min\{p_{1}(X),1-p_{0}(X)\}$, and the proportions of principal strata by $e_{s,\eta}=\mathbb{E}\{e_{s,\eta}(X)\}$ for all $s$. When $\eta(X)=1$, strata $10$ and $01$ have equal proportions, which corresponds to the situation where the treatment has zero average causal effect on $S$ and the principal scores are not identifiable (Jiang et al., 2022, Section S4.3). Thus, this boundary case was ruled out. The identification of the PGCEs without monotonicity also requires an alternative version of the principal ignorability (Assumption S1), which we provide in Section S7.1 of the Online Supplement.
With a fixed value of the sensitivity function $\eta(X)$, we derive bias-corrected multiply robust and debiased machine learning estimators for the PGCEs, leveraging the EIFs obtained without assuming monotonicity. The EIFs and the explicit forms of the bias-corrected estimators are given in Section S7.1 of the Online Supplement. Importantly, these multiply robust estimators remain triply robust under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$, as summarized in the following result (proofs provided in Section S7.2 of the Online Supplement).
Theorem 4 (Multiply robustness and efficiency of bias-corrected estimators)
Suppose that Assumptions 1 and S1 in Section S7.1 of the Online Supplement hold with a known $\eta(X)\neq 1$, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$. Then, for $s\in\{10,01,00,11\}$, the estimator $\widehat{\tau}_{h,\eta,\mathrm{tr}}^{s}$ is (a) triply robust such that it is consistent for $\tau_{h,\eta}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; and (b) asymptotically linear with influence function $\varphi_{h,\eta}^{s}$, thereby achieving the local semiparametric efficiency bound under $\mathcal{M}_{\pi+p+\mu}$.
@@ 5 Simulation studies
We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for the PGCEs, including the three plug-in moment estimators motivated by the nonparametric identification results in Section S2 of the Online Supplement, the triply robust estimators with parametric working models given in (13), and the debiased machine learning estimators given in (14), under different data-generating processes.
@@ 5.1 Probabilistic index with an intermediate variable
In the first simulation study, we consider the case of the probabilistic index estimand introduced in Examples 1 and 3, where the potential outcomes are continuous and drawn from a normal distribution. We first generate four pre-treatment covariates $X_{i}=(X_{i1},X_{i2},X_{i3},X_{i4})^{\prime}$, where $X_{i1},X_{i2},X_{i3}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$, and $X_{i4}\overset{\text{i.i.d.}}{\sim}\mathcal{B}(0.5)$. Then, we draw the treatment assignment from $Z_{i}|X_{i}\sim\mathcal{B}\{\mathrm{expit}(-X_{i1}+0.5X_{i2}-0.25X_{i3}-0.1X_{i4})\}$. Finally, the potential intermediate variable and the potential outcome are drawn from $D_{i}(Z_{i})|Z_{i},X_{i}\sim\mathcal{B}\{\mathrm{expit}(-1+2Z_{i}+X_{i1}-0.8X_{i2}+0.6X_{i3}-X_{i4})\}$, and $Y_{i}(Z_{i})|D_{i},Z_{i},X_{i}\sim\mathcal{N}(10+2Z_{i}-D_{i}+8X_{i1}+6X_{i2}+9X_{i3}+7X_{i4},1)$ for $i\in\{1,\ldots,n\}$, where, $\mathrm{expit}(x)=1/\{1+\exp(-x)\}$. We consider a binary intermediate variable ($D$ can be considered as the treatment receipt in the context of non-compliance) and assume the interest lies in estimating all three principal probabilistic index estimands ($\tau_{h}^{10}$, $\tau_{h}^{00}$ and $\tau_{h}^{11}$ with $h(u,v)=\mathbbm{1}(u\geq v)$). In addition, we create transformed covariates, $\widetilde{X}_{i}=(\widetilde{X}_{i1},\widetilde{X}_{i2},\widetilde{X}_{i3},\widetilde{X}_{i4})^{\prime}$, where
@@
to simulate scenarios where the nuisance functions are misspecified. Specifically, we fit the model using $X_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{X}_{i}$ otherwise. It is worth noting that, although in theory, machine learning methods are flexible with richer model spaces as $n$ increases, model misspecification can still be relevant to the debiased machine learning estimators, as their estimation accuracy in finite samples can be affected by the quality of the input covariate feature matrix (Cheng and Li, 2025).
@@
Figure 1 summarizes the results of the five estimators in scenarios where, at most, one of the three nuisance functions is misspecified. Plug-in moment estimators are consistent whenever all relevant nuisance functions are correctly specified. In contrast, the triply robust and debiased machine learning estimators are consistent across all four scenarios, regardless of whether all nuisance functions or subsets of nuisance functions are correctly specified. Results of each considered estimator in scenarios where at least two of the three nuisance functions are misspecified are presented in Web Figure S1 in Section S6 of the Online Supplement. As expected, the plug-in moment estimators are generally biased in those scenarios. In contrast, although biased, the triply robust and debiased machine learning estimators overall exhibit improved estimation accuracy. Finally, compared to the triply robust estimator, the debiased machine learning estimator yields a slight improvement, as it is generally less biased.
@@ 5.2 Win ratio with an intermediate variable
We further consider the case of estimating the win ratio estimands introduced in Examples 2 and 4, with the final potential outcomes being ordinal. We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study. However, for the potential outcome $Y_{i}(z)\in\{1,2,3\}$, we consider the following three-category ordinal logistic regression model:
@@
for $q\in\{1,2\}$, where $\zeta_{1,0}=-1$ and $\zeta_{2,0}=1$. Transformed covariates are created following Section 5.1. For the plug-in moment estimators and the triply robust estimator with parametric working models, the pairwise outcome mean is estimated using ordered logit regression, while treatment propensity and principal scores are estimated using logistic regression. For the debiased machine learning estimator, we used multinomial generalized linear models with the elastic net penalty from the glmnet package in R to estimate the pairwise outcome mean function. The treatment propensity and principal scores were estimated using the SuperLearner. The debiased machine learning estimator employs a five-fold cross-fitting procedure with $K=5$ and $L=15$. We set the number of observations $n=2,000$ with $1,000$ replications. In parallel to Section 5.1, violin plots summarizing the simulation results of each considered estimator are given in Figure 2 and Web Figure S2 in Section S6 of the Online Supplement.
From Figure 2, we observe that the plug-in moment estimators are consistent whenever all relevant nuisance functions are correctly specified. In contrast, the triply robust and debiased machine learning estimators are consistent in all four scenarios where at most one of the three nuisance functions is misspecified. When at least two of the three nuisance functions are misspecified (as shown in Web Figure S2 in Section S6 of the Online Supplement), all considered estimators are biased, with the triply robust and debiased machine learning estimators outperforming the others in terms of estimation accuracy (bias).
(a) $\tau_{h}^{10}$
(b) $\tau_{h}^{00}$
(c) $\tau_{h}^{11}$
Figure 2: Violin plots of estimators for the win ratio estimands in scenarios where at most one nuisance function is misspecified. Labels: ‘tp’ for treatment propensity or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment propensity (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (13); and ‘DML’ for the debiased machine learning estimators given in (14). Horizontal dashed lines represent the true values of the PGCEs.
@@ 5.3 Additional simulations under non-monotonicity
We consider a third simulation study to demonstrate the potential bias in the uncorrected triply robust and debiased machine learning estimators, as well as to validate the bias-corrected estimators under non-monotonicity. To achieve these goals, we modify the data-generating processes in Sections 5.1 and 5.2 by integrating the sensitivity function $\eta(X)$. Specifically, continuing with $p_{z}(X)=\mathrm{expit}(-1+2z+X_{1}-0.8X_{2}+0.6X_{3}-X_{4})$ for $z\in\{0,1\}$, we adopt the following configuration for the sensitivity function: $\eta(X)=\eta_{0}[1-\{p_{1}(X)-p_{0}(X)\}/\min\{p_{1}(X),1-p_{0}(X)\}]$, where, we let $\eta_{0}\in[0,1]$ to ensure the true sensitivity function is well-defined within the permissible bound. We consider two scenarios: one with moderate violations, represented by $\eta_{0}=0.5$, and another with severe violations, represented by $\eta_{0}=0.99$. The strata membership $S$ is drawn from from a multinominal distribution $\mathcal{M}\mathcal{N}\{e_{10,\eta}(X),e_{01,\eta}(X),e_{00,\eta}(X),e_{11,\eta}(X)\}$, where $\{e_{s,\eta}(X),s\in\{10,01,00,11\}\}$ are obtained following (15). The treatment assignment $Z\sim\mathcal{B}\{\mathrm{expit}(-X_{1}+0.5X_{2}-0.25X_{3}-0.1X_{4})\}$, and the intermediate variable $D$ is determined once $Z$ and $S$ are known. That is, $D=1$ if $Z=1,S\in\{10,11\}$ or $Z=0,S\in\{01,11\}$, and $D=0$ if $Z=1,S\in\{01,00\}$ or $Z=0,S\in\{10,00\}$. The potential outcome $Y(z)$ for $z\in\{0,1\}$ is generated following (16), based on the same covariates simulated as in Section 5.1.
To focus on the main ideas in this sensitivity simulation, we simplify without further considering nuisance function misspecification. We compare four estimators: the triply robust and debiased machine learning estimators under monotonicity, and the bias-corrected estimators with a known and correctly specified $\eta(X)$. The results are presented in Web Figure S3 in Section S6 of the Online Supplement.
Overall, the simulation results show that the bias-corrected estimators give consistent estimates of all four PGCEs under both non-monotonicity scenarios. As $\eta_{0}$ increases from $0.5$ to $0.99$, the biases of the triply robust and debiased machine learning estimators increase accordingly, due to a higher degree of non-monotonicity. Interestingly, such increases in bias are not uniform across all PGCEs. For example, the biases of $\widehat{\tau}_{h,\mathrm{tr}}^{10}$ and $\widehat{\tau}_{h,\mathrm{dml}}^{10}$ are relatively small under both $\eta_{0}$ values, suggesting that they may be dramatically affected by violations of the monotonicity assumption in our setup. On the other hand, the biases of $\widehat{\tau}_{h,\mathrm{tr}}^{00}$, $\widehat{\tau}_{h,\mathrm{dml}}^{00}$, $\widehat{\tau}_{h,\mathrm{tr}}^{11}$, and $\widehat{\tau}_{h,\mathrm{dml}}^{11}$ increased significantly, indicating their relatively higher sensitivity to non-monotonicity. In further exploration, we find that as $\eta_{0}$ increases, a larger proportion of the $01$ stratum comes from the $00$ and $11$ strata, hence contributing to increased bias in the estimators for $\tau_{h}^{00}$ and $\tau_{h}^{11}$.
@@ 6 Data example
We illustrate the proposed methods with a data example from the U.S. Jobs Corps study (Schochet et al., 2001), where individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged from 16 to 24. The data were from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either the treatment group, which received an immediate offer to participate in Job Corps, or the control group, which did not participate in Job Corps until three years later. As in many open-label randomized controlled trials, substantial noncompliance with treatment assignment occurred as an intermediate outcome 1 year after randomization: 15.4% ($857/5,577$) of participants assigned to treatment did not adhere to the assignment, and 50.6% ($1,854/3,663$) assigned to control switched to the treatment condition.
The proposed methods were implemented to examine the causal effect of the Job Corps intervention, with weekly earnings one year after the assignment selected as the final outcome. As the earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrasts may provide complementary insights into the treatment effects. In each data analysis below, we adjust for baseline covariates, including demographics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.

### v5

@@ 4.4 Sensitivity analysis for monotonicity
To point identify PGCEs, a central assumption is monotonicity (Assumption 2), which rules out the $01$ stratum. This stratum may be relevant, for example, when comparing two active interventions with unknown survival benefits against non-mortality outcomes. We thus describe a sensitivity analysis strategy to evaluate the robustness of the proposed estimators in the event of departure from monotonicity. Following the approach in Ding2016, we define the sensitivity function $\eta(\bm{X})$ as the ratio between the proportion of strata $01$ and $10$ conditional on the covariates, i.e., $\eta(\bm{X})=\mathbb{P}(S=01|\bm{X})/\mathbb{P}(S=10|\bm{X})$, where $\eta(\bm{X})\in[0,\infty)$; the monotonicity holds if $\eta(\bm{X})=0$, but $\eta(\bm{X})>0$ implies the existence of the $01$ stratum. The PCGEs in this scenario are thus $\tau_{h,\eta}^{s}$ for $s\in\{10,01,00,11\}$. To proceed, we first notice that, for $\eta(\bm{X})\neq 1$, the principal scores can be obtained as:
$e_{10,\eta}(\bm{X})=\frac{p_{1}(\bm{X})-p_{0}(\bm{X})}{1-\eta(\bm{X})},~~e_{01,\eta}(\bm{X})=\frac{\eta(\bm{X})\{p_{1}(\bm{X})-p_{0}(\bm{X})\}}{1-\eta(\bm{X})},$ (15)
$e_{00,\eta}(\bm{X})=1-p_{0}(\bm{X})-\frac{p_{1}(\bm{X})-p_{0}(\bm{X})}{1-\eta(\bm{X})},~~e_{11,\eta}(\bm{X})=p_{1}(\bm{X})-\frac{p_{1}(\bm{X})-p_{0}(\bm{X})}{1-\eta(\bm{X})},$
with $0\leq\eta(\bm{X})\leq 1-\{p_{1}(\bm{X})-p_{0}(\bm{X})\}/\min\{p_{1}(\bm{X}),1-p_{0}(\bm{X})\}$, and the proportions of principal strata by $e_{s,\eta}=\mathbb{E}\{e_{s,\eta}(\bm{X})\}$ for all $s$. When $\eta(\bm{X})=1$, strata $10$ and $01$ have equal proportions, which corresponds to the situation where the treatment has zero average causal effect on $S$, and the principal scores are not identifiable (Jiang2022, Section S4.3). Thus, this boundary case was ruled out. The identification of the PGCEs without monotonicity also requires an alternative version of the principal ignorability (Assumption S1), which we provide in Section S7.1 of the Online Supplement.
With a fixed value of the sensitivity function $\eta(\bm{X})$, we derive bias-corrected multiply robust and debiased machine learning estimators for the PGCEs, leveraging the EIFs obtained without assuming monotonicity. The EIFs and the explicit forms of the bias-corrected estimators are given in Section S7.1 of the Online Supplement. Importantly, these multiply robust estimators remain triply robust under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$, as summarized in the Theorem S6 in Section S7.1 of the Online Supplement.
@@ 5.1 Probabilistic index with an intermediate variable
In the first simulation study, we focus on the probabilistic index estimand introduced in Examples 1 and 3, in which the potential outcomes are continuous and drawn from a normal distribution. We first generate four pre-treatment covariates $\bm{X}_{i}=(X_{i1},X_{i2},X_{i3},X_{i4})^{\top}$, where $X_{i1},X_{i2},X_{i3}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$, and $X_{i4}\overset{\text{i.i.d.}}{\sim}\mathcal{B}(0.5)$. Then, we draw the treatment assignment from $Z_{i}|\bm{X}_{i}\sim\mathcal{B}\{\mathrm{expit}(-X_{i1}+0.5X_{i2}-0.25X_{i3}-0.1X_{i4})\}$. Finally, the potential intermediate variable and the potential outcome are drawn from $D_{i}(Z_{i})|Z_{i},\bm{X}_{i}\sim\mathcal{B}\{\mathrm{expit}(-1+2Z_{i}+X_{i1}-0.8X_{i2}+0.6X_{i3}-X_{i4})\}$, and $Y_{i}(Z_{i})|D_{i},Z_{i},\bm{X}_{i}\sim\mathcal{N}(10+2Z_{i}-D_{i}+8X_{i1}+6X_{i2}+9X_{i3}+7X_{i4},1)$ for $i\in\{1,\ldots,n\}$, where, $\mathrm{expit}(x)=1/\{1+\exp(-x)\}$. We consider a binary intermediate variable ($D$ can be considered as the treatment receipt in the context of non-compliance) and assume the interest lies in estimating all three principal probabilistic index estimands ($\tau_{h}^{10}$, $\tau_{h}^{00}$ and $\tau_{h}^{11}$ with $h(u,v)=\mathbbm{1}(u\geq v)$). In addition, we create transformed covariates, $\widetilde{\bm{X}}_{i}=(\widetilde{X}_{i1},\widetilde{X}_{i2},\widetilde{X}_{i3},\widetilde{X}_{i4})^{\top}$, where
@@
are used to simulate scenarios where the nuisance functions are misspecified. We fit the model using $\bm{X}_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{\bm{X}}_{i}$ otherwise. It is worth noting that, although in theory, machine learning methods are flexible with richer model spaces as $n$ increases, model misspecification can still be relevant to the debiased machine learning estimators, as their estimation accuracy in finite samples can be affected by the quality of the input covariate matrix (cheng2025inverting).
@@
Figure 1 summarizes the results of the five estimators in scenarios where, at most, one of the three nuisance functions is misspecified. Plug-in moment estimators are consistent whenever all relevant nuisance functions are correctly specified. In contrast, the triply robust and debiased machine learning estimators are consistent across all four scenarios, regardless of whether all nuisance functions or subsets of nuisance functions are correctly specified. Results of each considered estimator in scenarios where at least two of the three nuisance functions are misspecified are presented in Web Figure S1 in Section S6 of the Online Supplement. As expected, the plug-in moment estimators are generally biased in those scenarios. In contrast, although biased, the triply robust and debiased machine learning estimators overall exhibit improved estimation accuracy. Finally, compared with the triply robust estimator, the debiased machine learning estimator yields a slight improvement because it is generally less biased.
@@ 5.2 Win ratio with an intermediate variable
We further consider the case of estimating the win ratio estimands introduced in Examples 2 and 4, with the final potential outcomes being ordinal. We generate covariates $\bm{X}_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study. However, for the potential outcome $Y_{i}(z)\in\{1,2,3\}$, we consider the following three-category ordinal logistic regression model:
@@
for $q\in\{1,2\}$, where $\zeta_{1,0}=-1$ and $\zeta_{2,0}=1$. Transformed covariates are created following Section 5.1. For the plug-in moment estimators and the triply robust estimator with parametric working models, the pairwise outcome mean is estimated using ordered logit regression, while treatment propensity and principal scores are estimated using logistic regression. For the debiased machine learning estimator, we used multinomial generalized linear models with the elastic net penalty from the glmnet package in R to estimate the pairwise outcome mean function. The treatment propensity and principal scores were estimated using the SuperLearner. The debiased machine learning estimator employs a five-fold cross-fitting procedure with $K=5$ and $L=15$. We set the number of observations $n=2,000$ with $1,000$ replications. In parallel to Section 5.1, violin plots summarizing the simulation results of each considered estimator are given in Web Figures S2 and S3 in Section S6 of the Online Supplement.
From Web Figure S2, we observe that the plug-in moment estimators are consistent whenever all relevant nuisance functions are correctly specified. In contrast, the triply robust and debiased machine learning estimators are consistent in all four scenarios where at most one of the three nuisance functions is misspecified. When at least two of the three nuisance functions are misspecified (Web Figure S3 in Section S6 of the Online Supplement), all considered estimators are biased, with the triply robust and debiased machine learning estimators outperforming the others in terms of estimation accuracy (bias).
@@ 5.3 Additional simulations under non-monotonicity
We conduct a third simulation study to demonstrate bias in the uncorrected triply robust and debiased machine learning estimators, and to validate the bias-corrected estimators under non-monotonicity. We modify the data-generating processes in Sections 5.1 and 5.2 by integrating the sensitivity function $\eta(\bm{X})$. The detailed data-generating models and numerical results are presented in Section S7.2 of the Online Supplement. Overall, the results confirm that the bias-corrected estimators give consistent estimates of all four PGCEs across scenarios, suggesting their usefulness when monotonicity cannot be assumed.
@@ 6 Data example
We illustrate the proposed methods with a data example from the U.S. Jobs Corps study (Schochet2001), in which individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged 16 to 24. The data were from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either the treatment group, which received an immediate offer to participate in Job Corps, or the control group, which did not participate in Job Corps until three years later. As in many open-label randomized controlled trials, substantial noncompliance with treatment assignment occurred as an intermediate outcome 1 year after randomization: 15.4% ($857/5,577$) of participants assigned to treatment did not adhere to the assignment, and 50.6% ($1,854/3,663$) assigned to control switched to the treatment condition.
The proposed methods were implemented to examine the causal effect of the Job Corps intervention, with weekly earnings one year after the assignment selected as the final outcome. As earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrast may provide complementary insights into the treatment effects. In each data analysis below, we adjust for baseline covariates, including demographics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.